OpenStudy (sarahc):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10 (1, –2, –3) (–4, –1, –2) (–3, –2, –3) (2, –2, –4)

OpenStudy (sarahc):

B? :/

OpenStudy (sarahc):

@jim_thompson5910

OpenStudy (carson889):

It is indeed B.

OpenStudy (carson889):

\[\left[\begin{matrix}1 & 4& 1 & -10 \\ 3 & -3 & 6 & -21\\ 1&2&2&-10\end{matrix}\right]\] Row 1: R_1 Row 2: R_2 Row 3: R_3 1. R_1: R_1 - R_3 = 0,2,-1,0 2. R_3: R_3 - R_1 = 1,0,3,-10 3. R_2: R_2 - 3R_3 = 0,-3,-3,9 4. R_2: R_2 / -3 = 0,1,1,-3 5. R_1: R_1 - 2R_2 = 0,0,-3,6 **6. R_1: R_1 / -3 = 0,0,1,-2 ** 7. R_2: R_2 - R_1 = 0,1,0,-1 ** 8. R_3: R_3 - 3R_1 = 1,0,0,-4 Therfore x_1 = -4, x_2 = -1, x_3 = -2